# Text-guided Controllable Mesh Refinement for Interactive 3D Modeling

### v1

@@ Abstract.
We propose a novel technique for adding geometric details to an input coarse 3D mesh guided by a text prompt. Our method is composed of three stages. First, we generate a single-view RGB image conditioned on the input coarse geometry and the input text prompt. This single-view image generation step allows the user to pre-visualize the result and offers stronger conditioning for subsequent multi-view generation. Second, we use our novel multi-view normal generation architecture to jointly generate six different views of the normal images. The joint view generation reduces inconsistencies and leads to sharper details. Third, we optimize our mesh with respect to all views and generate a fine, detailed geometry as output. The resulting method produces an output within seconds and offers explicit user control over the coarse structure, pose, and desired details of the resulting 3D mesh. Project page: https://text-mesh-refinement.github.io.
@@ 1. Introduction
We consider the problem of adding geometric details to 3D object meshes through text guidance. Text-to-3D generative modeling has been successfully applied to computer vision (Poole et al., 2023), computer graphics (Khalid et al., 2022) and geometry processing (Gao et al., 2023) applications. These methods focus on generating a 3D mesh directly from text (Poole et al., 2023; Wang et al., 2023) without providing users with the ability to control the coarse structure of the output shape, limiting their practical usefulness for artists who need careful control over the generation process. Other methods focus on generating textures for 3D meshes (Richardson et al., 2023), but they do not perform any changes in the geometry. While there are some approaches (Gao et al., 2023; Metzer et al., 2023) that provide control to users and are able to modify the geometry of a given shape, these methods are generally slow since they rely on costly score distillation sampling (Poole et al., 2023).
@@ 2. Related Work
Another line of works finetuned image diffusion models on large 3D object datasets, such as Objaverse (Deitke et al., 2023), to arrive at diffusion models capable of generating multiple images of the same 3D object viewed from different viewpoints. Such methods can produce multi-view images from a text prompt or a single-view image that could be generated by a text-to-image model. A 3D shape can then be reconstructed from the multi-view images, either via optimizing a 3D shape with differentiable rendering (Liu et al., 2023a; Shi et al., 2023, 2024; Long et al., 2024; Liu et al., 2024a), or via a feed-forward 3D shape reconstruction network (Liu et al., 2023b, 2024b; Li et al., 2023; Hong et al., 2023).
@@ 3.1. Single-View RGB Generation
The goal of the first stage of our approach is to generate an RGB image that will act as a guide for the rest of the mesh refinement process. This is necessary for two reasons. First, this image can be generated using large-scale text-to-image models that are general enough to work with various prompts and types of objects while still being capable of creating detailed imagery. Second, this initial procedure can act as a fast preview of the whole process – the user can quickly verify (in $\sim$3 seconds) if the image corresponds to the appearance they intend before continuing with the rest of the pipeline. The process starts by the user providing a starting viewpoint $\theta_{s}$, an initial mesh $M_{\text{coarse}}$ and the text prompt $P$. We use a standard depth-to-RGB ControlNet model $\mathcal{G}_{d}$ to create a color image $I_{s}$ that will be used in the rest of the process:
@@ 3.3. Mesh Refinement and Optimization
Magic3D 0.2327 5597
Latent-NeRF 0.2367 3697
Fantasia3D 0.2556 2971
Ours 0.2666 32
4. Experiments
In this section, we analyze the capabilities of our method and investigate how modifying various components impacts its overall performance. We start by presenting a quantitative and qualitative comparison with the current state-of-the-art in the text-guided mesh refinement task. Then, we show how the user can apply different guidance levels to trade off between detail generation and original shape preservation. Finally, we present a few additional results and applications.
@@ 4.1. Text-guided Mesh Refinement
We choose Latent-NeRF (Metzer et al., 2023), Magic3D (Lin et al., 2023) and Fantasia3D (Chen et al., 2023a) as the competing methods. Latent-NeRF introduces global shape guidance through a sketch shape and a loss between the underlying NeRF’s occupancy value and the winding number to the sketch shape’s surface. We use their official implementation and their default setting to generate results for comparison with our method. Fantasia3D has their shape-guided 3D generation procedure implemented by initializing their initial SDF representation with an input shape. However, they don’t enforce any control mechanism during the generation. As a result, their generation can diverge from the input shape’s outline. Magic3D introduced a two-stage method for text-to-3D generation. Their second stage refines the coarse shape generated from the first stage via direct mesh optimization with differentiable rendering and it is therefore comparable to our method. For a fair comparison, we remove the first stage and initialize the Magic3D’s second stage geometry with the guidance shape in a randomly initialized color field. For both Magic3D and Fantasia3D, we use the public implementation by threestudio (Guo et al., 2023). Notice that all these methods rely on the score distillation sampling which usually takes more than 30 minutes per generation while our method takes less than a minute. For a runtime comparison, please refer to Table 1.
@@
Input 3D mesh
Wonder3D
Ours
Perceptual study and qualitative results.
Input 3D mesh
Ours w/o $\mathcal{L}_{\nabla n}$
Ours
@@ 4.3. Application
We develop a demo that builds on the Gradio library to provide users with a user-friendly interface for 3D modeling applications. See the demo video at https://text-mesh-refinement.github.io.
@@ 5. Conclusions
Acknowledgments
This paper is funded in part by NSERC Discovery (RGPIN–2022–04680), the Ontario Early Research Award program, the Canada Research Chairs Program, a Sloan Research Fellowship, the DSI Catalyst Grant program and gifts by Adobe Inc.
We thank Chenxi Liu, Otman Benchekroun, Joonho Kim, Kinjal Parikh, Kevin Mu, Abhishek Madan, Gabrielle Browne, Wenzhi Guo, Aravind Ramakrishnan, Victor Rong, Yixin Chen, Lily Goli, Yuta Noma, Anagh Malik, Chang Yue, Chao-Te Chou, Dylan Turpin, Tsun-Yi Yang, Yuliang Zou, Chen Gao, Shih-Yang Su and Wendy Choi for helping with the perceptual study, and Ziyi Wu, Shih-Yang Su and Yuliang Zou for insightful conversations that inspired us in this work. We would like to acknowledge the Vector Institute for the computing support. We thank Xuan Dam, John Hancock and all the University of Toronto Department of Computer Science research, administrative and maintenance staff.

### v2

@@ Abstract.
We propose a novel technique for adding geometric details to an input coarse 3D mesh guided by a text prompt. Our method is composed of three stages. First, we generate a single-view RGB image conditioned on the input coarse geometry and the input text prompt. This single-view image generation step allows the user to pre-visualize the result and offers stronger conditioning for subsequent multi-view generation. Second, we use our novel multi-view normal generation architecture to jointly generate six different views of the normal images. The joint view generation reduces inconsistencies and leads to sharper details. Third, we optimize our mesh with respect to all views and generate a fine, detailed geometry as output. The resulting method produces an output within seconds and offers explicit user control over the coarse structure, pose, and desired details of the resulting 3D mesh.
Text-to-3D, Generative AI
submissionid: 624journalyear: 2024copyright: acmlicensedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687630isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Computer graphics
@@ 1. Introduction
We consider the problem of adding geometric details to 3D object meshes through text guidance. Text-to-3D generative modeling has been successfully applied to computer vision (Poole et al., 2023; Zhang et al., 2024), computer graphics (Khalid et al., 2022) and geometry processing (Gao et al., 2023; Xu et al., 2024) applications. These methods focus on generating a 3D mesh directly from text (Poole et al., 2023; Wang et al., 2023) without providing users with the ability to control the coarse structure of the output shape, limiting their practical usefulness for artists who need careful control over the generation process. Other methods focus on generating textures for 3D meshes (Richardson et al., 2023; Cao et al., 2023), but they do not perform any changes in the geometry. While there are some approaches (Gao et al., 2023; Metzer et al., 2023) that provide control to users and are able to modify the geometry of a given shape, these methods are generally slow since they rely on costly score distillation sampling (Poole et al., 2023).
@@ 2. Related Work
Another line of works finetuned image diffusion models on large 3D object datasets, such as Objaverse (Deitke et al., 2023), to arrive at diffusion models capable of generating multiple images of the same 3D object viewed from different viewpoints. Such methods can produce multi-view images from a text prompt or a single-view image that could be generated by a text-to-image model. These multi-view images can be used to texture 3D meshes (Richardson et al., 2023) or can be used to reconstruct a 3D shape via optimizing a 3D shape with differentiable rendering (Liu et al., 2023a; Shi et al., 2023, 2024; Long et al., 2024; Liu et al., 2024a) or via a feed-forward 3D shape reconstruction network (Liu et al., 2023b, 2024b; Li et al., 2023; Hong et al., 2023).
@@ 3.1. Single-View RGB Generation
The goal of the first stage of our approach is to generate an RGB image that will act as a guide for the rest of the mesh refinement process. This is necessary for two reasons. First, this image can be generated using large-scale text-to-image models that are general enough to work with various prompts and types of objects while still being capable of creating detailed imagery. Second, this initial procedure can act as a fast preview of the whole process – the user can quickly verify (in $\sim$3 seconds) if the image corresponds to the appearance they intend before continuing with the rest of the pipeline. The process starts with the user providing a starting viewpoint $\theta_{s}$ (the view is given by the artist, i.e., the view that the object faces the camera), an initial mesh $M_{\text{coarse}}$ and the text prompt $P$. We use a standard depth-to-RGB ControlNet model $\mathcal{G}_{d}$ to create a color image $I_{s}$ that will be used in the rest of the process:
@@ 3.3. Mesh Refinement and Optimization
4. Experiments
In this section, we analyze the capabilities of our method and investigate how modifying various components impacts its overall performance. We start by presenting a quantitative and qualitative comparison with the current state-of-the-art in the text-guided mesh refinement task. Then, we show how the user can apply different guidance levels to trade off between detail generation and original shape preservation. Finally, we present a few additional results and applications. We will make our code and trained model weights available to the public upon acceptance.
@@ 4. Experiments
Magic3D 0.2327 $\pm$ 0.0393 5597
Latent-NeRF 0.2367 $\pm$ 0.0415 3697
Fantasia3D 0.2556 $\pm$ 0.0335 2971
Ours 0.2666 $\pm$ 0.0317 32
@@ 4.1. Text-guided Mesh Refinement
While there are methods that rely on supervisory signals from joint text-image embeddings (Gao et al., 2023; Khalid et al., 2022), recent work has shown that better reconstruction quality comes from methods using score distillation sampling (SDS) (Poole et al., 2023). For this reason, we mainly compare with methods using SDS. We choose Latent-NeRF (Metzer et al., 2023), Magic3D (Lin et al., 2023) and Fantasia3D (Chen et al., 2023a) as the competing methods. Latent-NeRF introduces global shape guidance through a sketch shape and a loss between the underlying NeRF’s occupancy value and the winding number to the sketch shape’s surface. We use their official implementation and their default setting to generate results for comparison with our method. Fantasia3D has their shape-guided 3D generation procedure implemented by initializing their initial SDF representation with an input shape. However, they don’t enforce any control mechanism during the generation. As a result, their generation can diverge from the input shape’s outline. Magic3D introduced a two-stage method for text-to-3D generation. Their second stage refines the coarse shape generated from the first stage via direct mesh optimization with differentiable rendering and it is therefore comparable to our method. For a fair comparison, we remove the first stage and initialize the Magic3D’s second stage geometry with the guidance shape in a randomly initialized color field. For both Magic3D and Fantasia3D, we use the public implementation by threestudio (Guo et al., 2023). Notice that all these methods rely on the score distillation sampling which usually takes more than 30 minutes per generation while our method takes less than a minute. For a runtime comparison, please refer to Table 1.
Input 3D mesh
Wonder3D
Ours
@@ 4.3. Application
We develop a demo that builds on the Gradio library to provide users with a user-friendly interface for 3D modeling applications. See the supplemental video for the demo. We will release our demo upon paper acceptance.
Input 3D mesh
Ours w/o $\mathcal{L}_{\nabla n}$
Ours
@@ 5. Conclusions
Acknowledgements.
This paper is funded in part by NSERC Discovery (RGPIN–2022–04680), the Ontario Early Research Award program, the Canada Research Chairs Program, a Sloan Research Fellowship, the DSI Catalyst Grant program and gifts by Adobe Inc. We thank Chenxi Liu, Otman Benchekroun, Joonho Kim, Kinjal Parikh, Kevin Mu, Abhishek Madan, Gabrielle Browne, Wenzhi Guo, Aravind Ramakrishnan, Victor Rong, Yixin Chen, Lily Goli, Yuta Noma, Anagh Malik, Chang Yue, Chao-Te Chou, Dylan Turpin, Tsun-Yi Yang, Yuliang Zou, Chen Gao, Shih-Yang Su and Wendy Choi for helping with the perceptual study, and Sven Dickinson, Sanja Fidler, Ziyi Wu, Shih-Yang Su and Yuliang Zou for insightful conversations that inspired us in this work. We would like to acknowledge the Vector Institute for the computing support. We thank Xuan Dam, John Hancock and all the University of Toronto Department of Computer Science research, administrative and maintenance staff.